A Spanish class has a total of 40 students. The number of males is 4 more than the number of females. How many males and how many females are in the class?
Answer by fcabanski(1391)

f+m=40 so m = 40 - f.


m = f+4 (The number of males is four more than the number of females.)


40-f = f+4 ---> 36 = 2f ---> f = 18.


m = f+ 4 = 18+4 = 22

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)